Bette’s Bananas delivers fresh bananas to your doorstep every morning. The cost for the service is $12/month flat rate and your first ten bananas are free! Further bananas cost $0.10 each. Construct a piecewise-linear function for the cost C of receiving N bananas in a month.
Answer by uma(370)


Let the number of bananas received be = N
Number of bananas free = 10
==> number of bananas to be charged = N-10
Cost of N-10 bananas at $0.10 each = (N-10)* 0.10
Cost for the service per month = $12
Therefore the cost C for receiving N bananas is given by,
C = 12 + (N-10)*(0.10)

Good Luck!!!